OpenStudy (anonymous):

Three waiters at a local restaurant earned different amounts of money the past week. Jennifer earned $40 less than twice the amount Tony earned. Chris earned $60 more than Tony. How much did each waiter earn if the total of all three was $900?

OpenStudy (anonymous):

900 = t (tony) + 2t-40 (jennifer) + t+60 (chris) simplify 900 = 4t+20 880 = 4t simplify 220 = t Tony = $220 Chris = 60+220 = $280 Jennifer = 2t-40 = $400 Check work: 400+280+220= $900